### v4

@@ 4.2 Avoidance of strictly saddle points
In this subsection, we further show that the DIRL2 algorithm does not converge to fixed points that are stationary points and strict saddle points. Unlike DIRL1, the map $S$ in DIRL2 does not enjoy the property of “active manifold identification” or “model identification”. There is no guarantee that the iteration map is $C^{1}$-smooth without this property. The following additional assumption is needed to guarantee the local smoothness. One can verify that most sparse nonconvex penalty in Table 1 like LPN, TAN, FRA, EXP and LOG are satisfied.
@@
$\lim\limits_{z\rightarrow 0}r^{\prime}(|z|)=\infty\ \text{ and }\ \lim\limits_{z\rightarrow 0}z\frac{r^{\prime\prime}(|z|)}{r^{\prime}(|z|)^{2}}\to 0.$
To show the smoothness of $S$, we first calculate the Jacobian of $S$. Define
$g(z):=\frac{1}{1+\tfrac{\lambda}{\beta}\tfrac{r^{\prime}(|z|)}{|z|}}\ \text{ with }g(0):=0.$
@@
$g^{\prime}(z)=\begin{cases}-\tfrac{\lambda}{\beta}\frac{r^{\prime\prime}(|z|)z-\text{sign}(z)r^{\prime}(|z|)}{(|z|+\tfrac{\lambda}{\beta}{r^{\prime}(|z|)})^{2}},&\text{\ for }z\neq 0,\\ \lim\limits_{\delta\to 0}\left(\frac{1}{1+\tfrac{\lambda}{\beta}\tfrac{r^{\prime}(|\delta|)}{|\delta|}}-0\right)\frac{1}{\delta}=0,&\text{\ for }z=0.\end{cases}$ (39)
@@
$\lim\limits_{z\to 0}g^{\prime}(z)=$ $-\tfrac{\lambda}{\beta}\lim\limits_{z\to 0}\left(\frac{r^{\prime\prime}(|z|)z}{(|z|+\tfrac{\lambda}{\beta}{r^{\prime}(|z|)})^{2}}-\frac{\text{sign}(z)r^{\prime}(|z|)}{(|z|+\tfrac{\lambda}{\beta}{r^{\prime}(|z|)})^{2}}\right)$
$=$ $-\lim\limits_{z\to 0}\left(\tfrac{\beta}{\lambda}\frac{r^{\prime\prime}(|z|)z}{{r^{\prime}(|z|)}^{2}}-\text{sign}(z)\frac{|z|}{r^{\prime}(|z|)}\right)=0=g^{\prime}(0)$
@@
We have the following result for $S$ and $T$.
Theorem 19.
(Local smoothness and Lipschitz continuity) Suppose Assumption 1-4 hold. The following statements hold:
@@
    $DS(x^{*},\bm{0})=\ \begin{bmatrix}H_{\mathcal{I}^{*}\mathcal{I}^{*}}&H_{\mathcal{I}^{*}\mathcal{J}^{*}}&,\bm{0}\\ 0&0&0\\ 0&0&\mu I\end{bmatrix},$ (40)
    where $H_{\mathcal{I}^{*}\mathcal{I}^{*}}=I-\tfrac{1}{\beta}(I+\tfrac{\lambda}{\beta}W_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$.
@@
    $DT(x^{*},\bm{0})=\ \begin{bmatrix}(1-\alpha)I+\alpha H_{\mathcal{I}^{*}\mathcal{I}^{*}}&H_{\mathcal{I}^{*}\mathcal{J}^{*}}&\bm{0}\\ \bm{0}&(1-\alpha)I&\bm{0}\\ \bm{0}&\bm{0}&(1-\alpha(1-\mu))I\end{bmatrix}.$
Proof.
(i) We simply calculate its partial derivative and verify that it is continuous at any point $(x^{*},\epsilon^{*})\in\mathbb{R}^{2n}$.
@@
$\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ g(z^{*}_{i})(1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*}))+\tfrac{x_{i}^{*}}{z_{i}^{*}}g^{\prime}(z_{i}^{*})(x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*}));$ (41)
$\frac{\partial}{\partial x_{j}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ -g(z^{*}_{i})\tfrac{1}{\beta}\nabla^{2}_{ij}f(x^{*}),\quad j\neq i;$ (42)
$\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x^{*},\epsilon^{*})=$ $\ \tfrac{\epsilon_{i}^{*}}{z_{i}^{*}}g^{\prime}(z^{*}_{i})(x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*}));$ (43)
$\frac{\partial}{\partial\epsilon_{j}}S_{i}^{\epsilon}(x^{*},\epsilon^{*})=$ $\ 0,\quad j\neq i.$ (44)
Consider $i$ satisfying $x_{i}^{*}=0$ and $\epsilon_{i}^{*}=0$. It follows from $S_{i}^{x}(x^{*},\bm{0})=0$ and $r^{\prime}(|0^{+}|)=\infty$ that
$\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ \lim\limits_{\delta\rightarrow 0}\frac{1}{1+\frac{2\lambda}{\beta}w(\delta,0)}[\delta-\tfrac{1}{\beta}\nabla_{i}f(x^{*}+\delta e_{i})]/\delta$ (45)
@@
$\frac{\partial}{\partial x_{j}}S_{i}(x^{*},\epsilon^{*})=$ $\ \lim\limits_{\delta\rightarrow 0}\frac{1}{1+\frac{2\lambda}{\beta}w(\delta,0)}[\delta-\tfrac{1}{\beta}\nabla_{i}f(x^{*}+\delta e_{j})]/\delta$ (46)
@@
$\frac{\partial}{\partial\epsilon_{i}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ \lim\limits_{\delta\rightarrow 0}\frac{1}{1+\frac{2\lambda}{\beta}w(0,\delta)}[-\tfrac{1}{\beta}\nabla_{i}f(x^{*})]/\delta$ (47)
@@
Now we verify that the Jacobian is continuous at $(x^{*},\epsilon^{*})$. It suffices to show this for any $i$ satisfying $x_{i}^{*}=0$ and $\epsilon_{i}^{*}=0$; indeed, it holds that It follows from (41), the fact that $\frac{x_{i}}{z_{i}}\leq 1$ and $g(0)=0,g^{\prime}(0)=0$ that
$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial x_{i}}S_{i}^{x}(x,\epsilon)=0=\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*}).$
@@
$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial x_{j}}S_{i}^{x}(x,\epsilon)=0=\frac{\partial}{\partial x_{j}}S_{i}^{x}(x^{*},\epsilon^{*}).$
@@
$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x,\epsilon)=0=\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x^{*},\epsilon^{*}).$
@@
$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial\epsilon_{j}}S_{i}^{\epsilon}(x,\epsilon)=0=\frac{\partial}{\partial\epsilon_{j}}S_{i}^{\epsilon}(x^{*},\epsilon^{*}).$
@@
To derive the matrix form of the Jacobian, it follows from (41) that for $i\in\mathcal{I}^{*}$,
$h_{ii}:=$ $\ \frac{\partial}{\partial x_{j}}S_{i}(x^{*},\bm{0})$
@@
$=$ $\ \frac{(1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*}))}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}-\frac{\tfrac{\lambda}{\beta}\left(r^{\prime\prime}(|x^{*}_{i}|)-\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}\right)}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x_{i}^{*}|}}$
$=$ $\ \frac{1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*})-\tfrac{\lambda}{\beta}r^{\prime\prime}(|x_{i}^{*}|)+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}$
$=$ $\ 1-\tfrac{1}{\beta}\frac{\nabla^{2}_{ii}f(x^{*})+\lambda r^{\prime\prime}(|x_{i}^{*}|)}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}$
@@
$h_{ij}:=\frac{\partial}{\partial x_{j}}S_{i}(x^{*},\bm{0})=-\tfrac{1}{\beta}\frac{\nabla^{2}_{ij}f(x^{*})}{1+\tfrac{\lambda}{\beta}\tfrac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}.$
We then combine this with (43)–(48) to derive (40) of the Jacobian $DS(x^{*},\bm{0})$. The Lipschitz continuity of $S$ trivially leads to the boundedness of $\mathcal{L}(F^{0})$.
(ii) This is straightforward by applying [23, Lemma 2.13 and 2.14]. The Jacobian $DT(x^{*},\bm{0})$ is obtained directly by (i). ∎
@@
(Unstable Fixed Point) Suppose Assumption 1-4 hold. Let $x^{*}$ be a stationary point of (P). If $x^{*}$ is a strict saddle point of (P), then the corresponding $(x^{*},\epsilon^{*})$ with $\epsilon^{*}=\bm{0}$ is an unstable fixed point of $T$. Moreover, for $\alpha<\beta/\rho$, $DT(x^{*},\bm{0})$ is invertible.
Proof.
It follows from Theorem 19(ii) that we only have to check the eigenvalues of
$(1-\alpha)I+\alpha H_{\mathcal{I}^{*}\mathcal{I}^{*}}=I-\tfrac{\alpha}{\beta}(I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}),$ (49)
@@
where we have $W^{*}_{ii}=w(x_{i}^{*},0)<\infty,i\in\mathcal{I}^{*}$. If $\lambda_{\min}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))<0$, it follows from Lemma [49, Proposition 6.1] that $\lambda_{\min}((I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))<0$. Hence, $DT(x^{*},\bm{0})$ has an eigenvalue greater than 1, meaning $(x^{*},\bm{0})$ is the unstable fixed point of $T$. Moreover, since $\|\tfrac{\alpha}{\beta}(I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\|\tfrac{\alpha}{\beta}(I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\|\|\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\rho$, it is clear that for $\alpha<\beta/\rho$, $DT(x^{*},\bm{0})$ is invertible. ∎
Now we can derive the global escape property of ${\rm DIRL}_{2}$ algorithm.
Theorem 21.
@@
(Global Escape) Consider problem (P) with strict saddle property and the ${\rm DIRL}_{2}$ algorithm with Assumption 1-4. Let the condition of Proposition 19 holds and $\alpha<\beta/\rho$. Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by the Algorithm 10. Then, the set of initial points attracted by strict saddle points has zero Lebesgue measure.
Proof.
@@
It follows directly from Theorem 19 and 20 that: (i) map $T$ is local $C^{1}$-smooth around any stationary points; (ii) all stationary points $x^{*}$ are unstable fixed points; (iii) $T$ is a lipeomorphism. Then by Corollary 1, the set of initial conditions attracted by such fixed points has zero Lebesgue measure. ∎
We illustrate this property by applying DIRL2 to Example 1. We test $\alpha=0.1$ and $0.9$ with different initial points; first is $(3,3)$ representing an initial point far from the global minimizer and the second is $(0,0.043)$ very close to the saddle point. The trajectories are shown in Figure 4. In four cases, the algorithm always converges to the global minimizer $(0,1)$.
(a) $\alpha=0.9$ and initial point $(3,3)$.
(b) $\alpha=0.9$ and initial point $(0,0.043)$.
(c) $\alpha=0.1$ and initial point $(3,3)$.
(d) $\alpha=0.1$ and initial point $(0,0.043)$.
Figure 4: Trajectories of DIRL2.
Appendix
Lemma A.1.
Given $a,b,c\in\mathbb{R}^{n}_{+}$ with $a_{i}\leq b_{i}+c_{i}\quad\text{for all }i$. It holds that $\|a\|\leq\|b\|+\|c\|$.
Proof.
It follows that $a_{i}^{2}\leq b_{i}^{2}+2b_{i}c_{i}+c_{i}^{2}$, implying $\sum_{i=1}^{n}a_{i}^{2}\leq\sum_{i=1}^{n}(b_{i}^{2}+2b_{i}c_{i}+c_{i}^{2})$, or equivalently, $\|a\|^{2}\leq\|b\|^{2}+2b^{T}c+\|c\|^{2}$. By the Cauchy-Schwarz inequality, $2\sum_{i=1}^{n}b_{i}c_{i}\leq 2\|b\|\|c\|$. Therefore, $\|a\|^{2}\leq(\|b\|+\|c\|)^{2}$, completing the proof. ∎
Lemma A.2.
For $x,y\in\mathbb{R}$ and $\alpha\in(0,1]$, it holds that $x^{2}-((1-\alpha)x+\alpha y)^{2}\geq\alpha(x^{2}-y^{2})$.
Proof.
It follows that
$x^{2}-((1-\alpha)x+\alpha y)^{2}$ $=x^{2}-(1-\alpha)^{2}x^{2}-\alpha^{2}y^{2}-2\alpha(1-\alpha)xy$
$=(2\alpha-\alpha^{2})x^{2}-\alpha^{2}y^{2}-2\alpha(1-\alpha)xy$
$=\alpha(x^{2}-y^{2})+(\alpha-\alpha^{2})x^{2}+(\alpha-\alpha^{2})y^{2}-2(\alpha-\alpha^{2})xy$
$=\alpha(x^{2}-y^{2})+(\alpha-\alpha^{2})(x-y)^{2}$
$\geq\alpha(x^{2}-y^{2}),$
where the inequality is by $\alpha\in(0,1]$. ∎

### v5

@@ 4.2 Avoidance of strictly saddle points
In this subsection, we further show that the DIRL2 algorithm does not converge to fixed points that are stationary points and strict saddle points. Unlike DIRL1, the map $S$ in DIRL2 does not enjoy the property of “active manifold identification” or “model identification”. There is no guarantee that the iteration map is $C^{1}$-smooth without this property. The following additional assumption is needed to guarantee local smoothness, which is satisfied by the $\ell_{p}$ regularization.
@@
$\lim\limits_{z\rightarrow 0^{+}}r^{\prime}(z)=\infty\ \text{ and }\ \lim\limits_{z\rightarrow 0^{+}}z\frac{r^{\prime\prime}(z)}{r^{\prime}(z)^{2}}\to 0.$
To show the smoothness of $S$, we first calculate the Jacobian of $S$. Define the auxiliary function $g:\mathbb{R}_{+}\to\mathbb{R}$:
$g(z):=\frac{1}{1+\tfrac{\lambda}{\beta}\tfrac{r^{\prime}(z)}{z}}\ \text{ with }g(0):=0.$
@@
(39) $g^{\prime}(z)=\begin{cases}-\tfrac{\lambda}{\beta}\frac{r^{\prime\prime}(z)z-r^{\prime}(z)}{(z+\tfrac{\lambda}{\beta}{r^{\prime}(z)})^{2}},&\text{\ for }z\neq 0,\\ \lim\limits_{\delta\to 0}\left(\frac{1}{1+\tfrac{\lambda}{\beta}\tfrac{r^{\prime}(|\delta|)}{|\delta|}}-0\right)\frac{1}{\delta}=0,&\text{\ for }z=0.\end{cases}$
@@
$\lim\limits_{z\to 0}g^{\prime}(z)=$ $-\tfrac{\lambda}{\beta}\lim\limits_{z\to 0}\left(\frac{r^{\prime\prime}(z)z}{(z+\tfrac{\lambda}{\beta}{r^{\prime}(z)})^{2}}-\frac{r^{\prime}(z)}{(z+\tfrac{\lambda}{\beta}{r^{\prime}(z)})^{2}}\right)$
$=$ $-\lim\limits_{z\to 0}\left(\tfrac{\beta}{\lambda}\frac{r^{\prime\prime}(z)z}{{r^{\prime}(z)}^{2}}-\frac{z}{r^{\prime}(z)}\right)=0=g^{\prime}(0)$
@@
We have the following result for maps $S$ and $T$.
Theorem 4.8.
(Local smoothness and Lipschitz continuity) Suppose Assumptions 1-4 hold and $x^{*}$ is a stationary point of (P). Then the following statements hold:
@@
    (40) $DS(x^{*},\boldsymbol{0})=\ \begin{bmatrix}H_{\mathcal{I}^{*}\mathcal{I}^{*}}&H_{\mathcal{I}^{*}\mathcal{J}^{*}}&\boldsymbol{0}\\ \boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}\\ \boldsymbol{0}&\boldsymbol{0}&\mu I\end{bmatrix},$
    where $H_{\mathcal{I}^{*}\mathcal{I}^{*}}=I-\tfrac{1}{\beta}(I+\tfrac{\lambda}{\beta}U_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$.
@@
    $DT(x^{*},\boldsymbol{0})=\ \begin{bmatrix}(1-\alpha)I+\alpha H_{\mathcal{I}^{*}\mathcal{I}^{*}}&\alpha H_{\mathcal{I}^{*}\mathcal{J}^{*}}&\boldsymbol{0}\\ \boldsymbol{0}&(1-\alpha)I&\boldsymbol{0}\\ \boldsymbol{0}&\boldsymbol{0}&(1-\alpha(1-\mu))I\end{bmatrix}.$
Proof 4.9.
(i) We simply calculate its partial derivative and verify that it is continuous at any point $(x^{*};\epsilon^{*})\in\mathbb{R}^{2n}$. The smoothness of $S^{\epsilon}$ is obvious. We only verify the smoothness of $S^{x}$.
@@
(41) $\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ g(z^{*}_{i})(1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*}))+\tfrac{x_{i}^{*}}{z_{i}^{*}}g^{\prime}(z_{i}^{*})(x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*}));$
(42) $\frac{\partial}{\partial x_{j}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ -g(z^{*}_{i})\tfrac{1}{\beta}\nabla^{2}_{ij}f(x^{*}),\quad j\neq i;$
(43) $\frac{\partial}{\partial\epsilon_{i}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ \tfrac{\epsilon_{i}^{*}}{z_{i}^{*}}g^{\prime}(z^{*}_{i})(x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*}));$
(44) $\frac{\partial}{\partial\epsilon_{j}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ 0,\quad j\neq i.$
Consider $i$ satisfying $x_{i}^{*}=0$ and $\epsilon_{i}^{*}=0$. It follows from $S_{i}^{x}(x^{*},\boldsymbol{0})=0$ and $r^{\prime}(|0^{+}|)=\lim\limits_{\delta\to 0}u(\delta,0)=\lim\limits_{\delta\to 0}u(0,\delta)=\infty$ that
(45) $\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ \lim\limits_{\delta\rightarrow 0}\frac{1}{1+\frac{2\lambda}{\beta}u(\delta,0)}[\delta-\tfrac{1}{\beta}\nabla_{i}f(x^{*}+\delta e_{i})]/\delta$
@@
(46) $\frac{\partial}{\partial x_{j}}S^{x}_{i}(x^{*},\epsilon^{*})=$ $\ \lim\limits_{\delta\rightarrow 0}\frac{1}{1+\frac{2\lambda}{\beta}u(\delta,0)}[\delta-\tfrac{1}{\beta}\nabla_{i}f(x^{*}+\delta e_{j})]/\delta$
@@
(47) $\frac{\partial}{\partial\epsilon_{i}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ \lim\limits_{\delta\rightarrow 0}\frac{1}{1+\frac{2\lambda}{\beta}u(0,\delta)}[-\tfrac{1}{\beta}\nabla_{i}f(x^{*})]/\delta$
@@
Now we verify that the Jacobian is continuous at $(x^{*},\epsilon^{*})$. It suffices to show this for any $i$ satisfying $x_{i}^{*}=0$ and $\epsilon_{i}^{*}=0$. It follows from (41), the fact that $\frac{x_{i}}{z_{i}}\leq 1$ and $g(0)=0,g^{\prime}(0)=0$ that
$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial x_{i}}S_{i}^{x}(x,\epsilon)=0=\frac{\partial}{\partial x_{i}}S_{i}^{x}(0,0).$
@@
$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial x_{j}}S_{i}^{x}(x,\epsilon)=0=\frac{\partial}{\partial x_{j}}S_{i}^{x}(0,0).$
@@
$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x,\epsilon)=0=\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(0,0).$
@@
$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial\epsilon_{j}}S_{i}^{\epsilon}(x,\epsilon)=0=\frac{\partial}{\partial\epsilon_{j}}S_{i}^{\epsilon}(0,0).$
@@
To derive (40), now suppose $x^{*}$ is a stationary point for (P) and $\epsilon^{*}=\boldsymbol{0}$.
For $i\in\mathcal{I}^{*}$, (5) is equivalent to $x_{i}^{*}-\frac{1}{\beta}\nabla_{i}f(x^{*})=x_{i}^{*}+\frac{\lambda}{\beta}\frac{x_{i}^{*}}{z_{i}^{*}}r^{\prime}(|x^{*}_{i}|)$ and $x_{i}^{*}/z_{i}^{*}=\text{sign}(x_{i}^{*})$. It follows from (41) that for $i\in\mathcal{I}^{*}$,
(49) $h_{ii}=$ $\ \frac{\partial}{\partial x_{j}}S^{x}_{i}(x^{*},\boldsymbol{0})$
@@
$=$ $\ 1-\tfrac{1}{\beta}\frac{\nabla^{2}_{ii}f(x^{*})+\lambda r^{\prime\prime}(|x_{i}^{*}|)}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}.$
@@
(50) $h_{ij}:=\frac{\partial}{\partial x_{j}}S_{i}(x^{*},\boldsymbol{0})=-\tfrac{1}{\beta}\frac{\nabla^{2}_{ij}f(x^{*})}{1+\tfrac{\lambda}{\beta}\tfrac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}.$
We then combine this with (43)–(50) to derive (40). The boundedness of $\mathcal{L}(F^{0})$ trivially leads to the Lipschitz continuity of $S$.
(ii) This is straightforward by applying [15, Lemma 2.13 and 2.14]. The Jacobian $DT(x^{*},\boldsymbol{0})$ is obtained directly by (i).
@@
(Unstable Fixed Point) Suppose Assumptions 1-4 hold. If $x^{*}$ is a strict saddle point of (P), then the corresponding $(x^{*},\epsilon^{*})$ with $\epsilon^{*}=\boldsymbol{0}$ is an unstable fixed point of $T$. Moreover, for $\alpha<\beta/\rho$, $DT(x^{*},\boldsymbol{0})$ is invertible.
Proof 4.11.
By Theorem 4.8(ii), it suffices to check the eigenvalues of
(51) $(1-\alpha)I+\alpha H_{\mathcal{I}^{*}\mathcal{I}^{*}}=I-\tfrac{\alpha}{\beta}(I+\tfrac{\lambda}{\beta}U^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}),$
@@
where we have $U^{*}_{ii}=u(x_{i}^{*},0)<\infty,i\in\mathcal{I}^{*}$. If $\lambda_{\min}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))<0$, it follows from Lemma [36, Proposition 6.1] that $\lambda_{\min}((I+\tfrac{\lambda}{\beta}U^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))<0$. Hence, $DT(x^{*},\boldsymbol{0})$ has an eigenvalue greater than 1, meaning $(x^{*},\boldsymbol{0})$ is the unstable fixed point of $T$. Moreover, since $\|(I+\tfrac{\lambda}{\beta}U^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\|(I+\tfrac{\lambda}{\beta}U^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\|\|\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\|\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\rho$ by (22) and $\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$ does not have 0 eigenvalue by the strict saddle property, we know by (22) that $DT(x^{*},\epsilon^{*})$ is invertible for $\alpha<\beta/\rho$.
Now we can derive the global escape property for the ${\rm DIRL}_{2}$ algorithm.
Theorem 4.12.
@@
(Global Escape) Consider problem (P) with strict saddle property and the ${\rm DIRL}_{2}$ algorithm with Assumptions 1-4. Let $\alpha<\beta/\rho$ and $\alpha\in(0,\frac{1}{1+L_{S}})$ where $L_{S}$ is as defined in Theorem 4.8. Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by the Algorithm 2. Then, the set of initial points attracted by strict saddle points has zero Lebesgue measure.
Proof 4.13.
@@
It follows directly from Theorems 4.8 and 4.10 that: (i) map $T$ is local $C^{1}$-smooth around any stationary point; (ii) all stationary points $x^{*}$ are unstable fixed points; (iii) $T$ is a lipeomorphism. Then by Corollary 2.15, the set of initial conditions attracted by such fixed points has zero Lebesgue measure.
Appendix A Proof supplementary